Financial planner Charlie Labrador recommends to his client to invest $100,000 of her life savings in a bond mutual fund with an expected return of 5% and standard deviation of 15% and the remaining $300,000 of her savings in a stock mutual fund with an expected return of 10% and standard deviation of 35%. The correlation between the two funds is 0.4. What would be the standard deviation of her total portfolio?
Answer #1

Weight of Bond Mutual Fund =100000/(100000+300000) =0.25
Weight of Stock Mutual Fund =1-0.25 =0.75
Standard Deviation of Her total Portfolio =((Weight of Bond*Standard Deviation)^2+(Weight of Equity*Standard Deviation )^2+2*Weight of Bond*Standard Deviation*weight of Equity*Standard Deviation*Correlation))^0.5
=(0.25*15%)^2+(0.75*35%)^2+2*0.25*0.75*15%*35%*0.4)^0.5 =27.96%